OpenStudy (anonymous):

A jar contains $29\frac{5}{7}$ tablespoons of peanut butter. If one serving of peanut butter is 2 tablespoons, how many servings of peanut butter does the jar contain? Express your answer as a mixed number.

OpenStudy (aaronq):

\(\sf servings =\dfrac{total ~\#~of~tablespoons~in ~jar}{tablespoon ~per~serving}=\dfrac{29\dfrac{5}{7}}{2}\) \(\sf \dfrac{\dfrac{(29*7)+5}{7}}{2}=\dfrac{\dfrac{203+5}{7}}{2}\) \(\sf \dfrac{\dfrac{208}{7}}{2}=\dfrac{208}{7*2}=\dfrac{208}{14}=\dfrac{104}{7}=14\dfrac{6}{7} \)